D =200-25P. , S =-16+11P. Where D is demand in thousands and P is price in cedis. Assuming demand increases as result of an increase in disposable income consumers while supply remains the same.



b. What will be the effect on equilibrium quantity and equilibrium price

1
Expert's answer
2022-05-09T14:51:35-0400

"D=200-25P,"

"S=-16+11P,"

"D=S,"

"200-25P=-16+11P,"

"P=6,"

"D=S=50."